What are the concentrations of OH− and H+ (H+ is short for H3O+) in a 0.00070 M solution of Ba(OH)2 at 25 °C? Assume complete dissociation.


1
Expert's answer
2022-04-26T03:16:04-0400

Solution:

Ba(OH)2 is the strong base that completely dissociates into barium (Ba2+) and hydroxide (OH) ions.

The dissociation equation for Ba(OH)2 in aqueous solution is:

Ba(OH)2 → Ba2+ + 2OH

From the reaction, one barium ion and two hydroxide ions are formed.

So, the concentration of OH is two times the concentration of Ba(OH)2

Therefore,

[OH] = 2 × CM(Ba(OH)2) = 2 × (0.00070 M) = 0.0014 M

[OH] = 0.0014 M


Kw = [H+] × [OH] = 1.0×10−14

Therefore,

[H+] = Kw / [OH] = (1.0×10−14) / (0.0014) = 7.14×10−12

[H+] = 7.14×10−12 M


Answer:

[OH] = 0.0014 M

[H+] = 7.14×10−12 M